An archer shoots arrows at a circular target where the central portion of the target
inside is called the bull. The archer hits the bull with probability 1/32. Assume that
the archer shoots 96 arrows at the target, and that all shoots are independent. What
is an approximated probability that an archer hit not more than one bull?
1
Expert's answer
2020-06-08T19:22:02-0400

Here it is given that probability of hitting the bull is 1/32


And all events are independent.


And he shots 96 arrows at the bull and we need to determine the probability such that archer hit not more than one bull.


Probability = P(shooting 0 times out of 96) + P(shooting 1 time out of 96) = "(1- \frac{1}{32})^{96} + 96 \times (1- \frac{1}{32})^{95} \frac{1}{32} = 0.1944315"